OpenStudy (anonymous):

Two years ago a father was four times as old as his son three years from now he will be only three times as old as his son how old is each now?

OpenStudy (anonymous):

(f-2) = 4(s-2)--------------1 f+3 = 3(s+3)--------------2

OpenStudy (anonymous):

x father's age now y son's age now \[\begin{cases}(x-2)=4(y-2)\\(x+3)=3(y+3)\end{cases}\]

OpenStudy (anonymous):

1) f = 4s-8+2 = 4s-6 2) f = 3s+9 - 3 = 3s + 6 so, 3s+6 = 4s - 6 s=12 f = 42

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOU THANK YOU THANK YOU!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

welcome welcome welcome